In a completely randomized design, 12 experimental units were used for the first treatment, 15 for the second treatment, and 20 for the third treatment. Complete the following analysis of variance (to 2 decimals, if necessary). Round p-value to four decimal places. If your answer is zero enter "0".

Source of Variation Sum of Squares Degrees of Freedom Mean Square F p-value

Treatments 1,300 --?-- --?-- ----?----- ---?------ ----?----

Error --?--      --?-- --?-- ----?----- ---?------ ----?----

Total 2,000      --?-- --?-- ----?----- ---?------ ----?----

At a .05 level of significance, is there a significant difference between the treatments?

The p-value is ?

What is your conclusion?
Answer #1

a) ANOVA TAble

SV SS Df MSS F P value
Treatment 1300 3-1=2 1300/2 =650 650/15.91=40.85 0.0
Error 2000-1300 =700 47-3=44 700/44=15.91
Total 2000 47-1=46

Since p value is less than the level of significance , we reject null hypothesis and conclude that there is significant difference between means of three treatments